Corn, one of the major “energy crops”, can produce about 2000 liter of Ethanol biofuel per acre per year. If your car need about 100 liter of ethanol fuel per week for your work commute, about how much land do you need to grow your own fuel?

Select one:

a. 0.5 acre

b. 20 acre

c. 2.6 acre

d. 25 acre
Answer #1

Here,

for the total amount of fuel needed

total amount of fuel = 100 * 52 = 5200 liter

total area needed = 5200/2000

total area needed = 2.6 acre

the total area is c) 2.6 acre